OpenStudy (anonymous):

Find the y-intercept for 5x + 8y = -24

OpenStudy (bibby):

to find the y intercept, set x equal to 0 and solve http://openstudy.com/code-of-conduct @tkhunny

OpenStudy (tkhunny):

@bibby Demonstrating the Intercept Form is an acceptable response. It is also acceptable to assist a student in obtaining tools, then watching the student progress over the course of a few questions. @harleeekoon You could try the Intercept form, as previously demonstrated.

OpenStudy (anonymous):

the formula for the y intercept is mx+b so what you have to do is first get y on one side to do that you could subtract 5x from one side as so: 5x+8y=-24 -5x -5x then you would have 8y=-5x-24 then to get rid of the 8 you would divide each side by 8 as so: 8y(underline then put 8 on bottom) then underline the -5x-24 then put 8 on the bottom. the 8s cancel out so your left with y= -5x-24 8 then simply divide the -24 by 8 which is 3 so the slope intercept is y=-5x-3 ~hope i helped sarah